OpenStudy (anonymous):

find deriv of f(x) =cot^3 (sq root x) plz show work and ty dont wanna fail my exam tmmrw D:

zepdrix (zepdrix):

\[\huge f(x)=\cot^3 \sqrt x\]Like this? :o

zepdrix (zepdrix):

So we'll work from the outside inward. We differentiate the outermost function first. In this case, the outermost function is the CUBE power, After we differentiate the outside, we make a copy of the inside, and multiply by it's derivative. \[\huge f'(x)=3\cot^2 \sqrt x(\color{royalblue}{\cot \sqrt x})'\] The prime is to show that we still need to take a derivative of this part.

zepdrix (zepdrix):

Taking the derivative of the blue part gives us,\[\huge f'(x)=3\cot^2 \sqrt x(\color{royalblue}{-\csc^2 \sqrt x})(\color{forestgreen}{\sqrt x})'\]

zepdrix (zepdrix):

We apply the chain rule again, multiplying by the derivative of the inner inner function, sqrt x.

zepdrix (zepdrix):

Taking the derivative of the green part gives us,\[\huge f'(x)=3\cot^2 \sqrt x(\color{royalblue}{-\csc^2 \sqrt x})\left(\color{forestgreen}{\frac{1}{2\sqrt x}}\right)\]

zepdrix (zepdrix):

Confused on any part of that? I know it's a ton of steps. If you can understand the chain rule ok, then it's a little easier to get through.